OpenStudy (anonymous):

The slope of the line whose equation is 3x - 2y = 4 is 2/3 3/2 -2

OpenStudy (anonymous):

ax + by + c = 0 Slope = -a/b

OpenStudy (anonymous):

so wat do u think in 3x - 2y - 4 =0

OpenStudy (anonymous):

m = -3 / -2 = 3/2

OpenStudy (anonymous):

i dont know what do you think it is

OpenStudy (anonymous):

Hope u got it nw